OpenStudy (anonymous):

someone please help me (√2- √−2)(√8 + √-2 )

OpenStudy (tkhunny):

Have you considered multiplication?

OpenStudy (anonymous):

\[\sqrt{-2}*\sqrt{-2}=\iota \sqrt{2}*\iota \sqrt{2}=i^2*\left( \sqrt{2} \right)^2=-\sqrt{2}\]

OpenStudy (anonymous):

correction not \[-\sqrt{2}\] it is -2

OpenStudy (anonymous):

\[\left( \sqrt{2}-\sqrt{-2} \right)\left( \sqrt{8}+\sqrt{-2} \right)=\sqrt{2*8}+\sqrt{2*-2}-\sqrt{-2*8}-2\] \[=4+2 \iota-4\iota-2=2-2\iota \]